# What computer science has to say about the simulation hypothesis

### v4

@@ 6.2 Running (self-)simulation using fully homomorphic encryption
Similar consequences arise if we consider self-simulation. We might be simulating ourselves, but have done so with an FHE version of ourselves. Similarly to the possibility discussed above, suppose this is the case, but that we have misplaced the decryption key. In this scenario, we are a simulation that we ourselves are running — but we cannot understand that simulation of ourselves, a simulation which is us.
@@ 7 The simulation graph
Section 6 contains a brief discussion of the philosophical issues that arise if we consider simulation involving more than two universes. Such situations are barely even considered in a cursory way in the existing literature. Yet, there are nontrivial mathematical properties of such situations. In particular, the graphical structure of universes simulating universes can be quite interesting.
@@ 7.1 The graph of simulations and self-simulations
Indeed, suppose we restrict attention to such a simulation graph universes where there is a directed edge from any node $V$ in the graph to any other node $V^{\prime}\neq V$. Then the simulation relation cannot be a partially ordered set. Given this, suppose that all of the nodes (universes) have a finite $W$ and an evolution function that obeys the PCT. As pointed out in Section 2.4, the set of such universes is countably infinite, and so we cannot assign a uniform probability distribution to that set. However, since the elements of the set are not partially ordered in the simulation graph, we also cannot assign a Cantor measure over the elements of the set, if we wish to use the simulation relation to fix how to assign such a measure to the nodes in the graph. This establishes the claim made in the introduction, that it is not possible to use a Cantor measure to assign probabilities to a very naturally defined set of universes (at least, it’s not possible if we try to use the simulation relation to fix the measure).
@@
A bare simulation graph could contain computational universes that exist in the same cosmological universe (see Section 2.2), obeying the same laws of physics, the same initial conditions, etc. It might also contain computational universes that exist in different cosmological universes. These two cases can be intermingled as well.
@@
Finally, suppose we restrict the simulation graph $\Gamma$ to only contain universes that can simulate themselves, Suppose as well that we restrict the graph so that there is at most a single edge from any universe to $V$ to $V^{\prime}$ (where $V^{\prime}$ may or may not differ from $V$). Then the directed edges in $\Gamma$ form a reflexive, transitive relation, i.e., they form a preorder. In general, the relation provided by the edges in that simulation graph can include both pairs of nodes that are symmetric under the relation and pairs of nodes that are anti-symmetric under the relation. So while it is a preorder, those edges need not provide either an equivalence relation or a partially ordered set.
@@
7.2 Some types of simulation graphs
There are many specific types of simulation graphs that might be interesting to explore. As an example, any particular universe $V$ might be running more than one simulation at once. Physically, this could occur by having one computer in that universe running multiple simulations simultaneously, just like a laptop runs multiple tasks simultaneously, by “swapping”. It could also occur by having multiple computers in the universe, all running simulations. [^20] In terms of the formalism in this paper, the latter case would mean that “the” computer $N$ of the universe is actually a set of multiple computers running independently, in parallel. The simulation lemma would apply directly, and the self-simulation lemma would also hold, where ${\mathcal{S}}(\Delta t)$ is the initial joint state of all of the computers in the universe. In such a situation, the fan-out from the node of that universe $V$ in the simulation graph would be greater than $1$.
There are several interesting special case of this example, where the universe $V$ is running simulations of multiple different universes, $V_{1},V_{2},\ldots$ simultaneously. Suppose that each of those different universes are running their own simulations, which in turn are running their own simulations, etc. In this case there is an extensive a branching structure of simulation paths coming out of $V$. Moreover, it could be that as one runs down some of those simulation paths they eventually converge on our universe. In this special case, we are simultaneously being simulated in more than one computer.
As another interesting special case, there might be beings in a universe who are running us in their computer, and who themselves are a simulation in a computer that we are running. In such a case, the aliens would be a simulation running in a computer that they themselves control, just “one step removed”. In such a situation, we too would be self-simulating “one step removed”. We would be a means for the aliens to “split” their ontological status in two, while also being a means for us to split our ontological status in two.
Taking this further, suppose we are being simulated by multiple other sets of aliens, but in addition we are running simulations that are us. We might even be running simulations of those beings who are running simulations of us. In the associated simulation graph there would be edges that are loops from us into us, and edges from those other universes into us, perhaps together with some edges from us into those other universes.
As a final comment, recall from Section 4.2 that when the conditions in Lemma 2 hold for a universe $V$, there are an infinite number of initial states of the computer, $n_{0}$, such that that computer simulates what the full universe $V$, including itself, will be. That infinite degeneracy may have some interesting philosophical and mathematical implications. For example, suppose I run a computer which is simulating what my entire universe would some time $\Delta t$ in my future, if my universe had my actual environment $w_{0}$ at $t=0$. Suppose though that that simulation is for an initial state of the computer, $n^{\prime}_{0}$, that is computationally equivalent to $n_{0}$, the actual simulation program I am using in that simulation — but where that program $n^{\prime}_{0}$ actually differs from $n_{0}$. In a certain sense, this would not be a perfect self-simulation, even though it would obey the self-simulation lemma. This possibility suggests replacing any single loop in the simulation graph (i.e., any single edge from a universe-node into itself) with a set of multiple such loops, distinguished from one another by the fact that they use different programs, all of which happen to be computationally equivalent.
@@ 7.3 Time-ordered and time-bounded simulation graphs
In general, a universe might only obey the RPCT partially, in that it can only simulate a subset $K\subset{\mathbb{N}}$ of all TMs $T^{k}$. A discussion of this case is presented just below, in Section 7.4.
For the moment though, consider a set of computable universes all of which obey the RPCT in full, for $K={\mathbb{N}}$. The simulation graph for such a set of universes is trivial; it is a fully connected graph. However, even if we’re considering a set of computable universes all of which obey RPCT for $K={\mathbb{N}}$, there might still be nontrivial structure in the graph given by placing an edge from $V$ to $V^{\prime}$ only if $V$ simulates $V^{\prime}$ sufficiently quickly. More formally, we can consider the time-bounded simulation graph in which there is an edge from $V$ to $V^{\prime}$ iff $V$ simulates $V^{\prime}$ with a function ${\mathcal{T}}(\Delta t^{\prime},w^{\prime},n^{\prime})$ that obeys some bound in the worst-case over all pairs $(w^{\prime},n^{\prime})$ in how fast it can grow as a function of $\Delta t^{\prime}$. For example, one could consider the variant of a simulation graph given by only placing an edge from $V$ to $V^{\prime}$ if for all pairs $(w^{\prime},n^{\prime})$, ${\mathcal{T}}(\Delta t^{\prime},w^{\prime},n^{\prime})$ grows at most polynomially with $\Delta t^{\prime}$. Other kinds of nontrivial simulation graphs arise if we consider the scaling of the resources (time, memory, etc.) needed to compute the functions ${\mathcal{T}},{\mathcal{W}},{\mathcal{N}}$, or in the case of self-simulation, $\mathcal{S}$}. (See also the discussion in Section 10 of time-minimal simulation functions, their scaling properties, and computational complexity theory.)
@@ 7.4 Weak RPCT
In general we do not need to assume the full strength of the RPCT to prove a particular instance of either the simulation lemma or the self-simulation lemma. In the case of the former, we just need to assume that the simulating universe can implement the evolution function of the universe being simulated, i.e., can implement the particular TM specified by that evolution function of the universe being simulated. It is not necessary that it can implement the Turing machine of any universe there is. In the case of the latter, we just need to assume that the simulating universe can implement the (Turing machine specifying the) computable function ${\mathcal{S}}(\Delta t)$. We do not need the computer in the simulating universe to be computationally universal.
Accordingly, given any $K\subset{\mathbb{N}}$, I say that the weak RPCT (for a set $K\subset{\mathbb{N}}$) holds for universe $V$ if Definition 4 holds for $V$ after rather than the requirement in Definition 4 that the functions $\widehat{{\mathcal{T}}},\widehat{{\mathcal{W}}}$ and $\widehat{{\mathcal{N}}}$ have the RPCT properties for all $k$, we instead only require they those three functions have the RPCT for all $k\in K$ (and evolution function $g$ of $V$).
The associated weak simulation lemma says that $V$ can simulate a universe $V^{\prime}=W^{\prime}\times N^{\prime}$ whose evolution function is given by the TM $T^{k}$ if
  1. 1.
    $V^{\prime}$ obeys the PCT;
  2. 2.
    $V$ obeys the weak RPCT for a set $K$ that includes three functions $\widehat{{\mathcal{T}}},\widehat{{\mathcal{W}}}$ and $\widehat{{\mathcal{N}}}$ that have the RPCT properties for all $(\Delta t^{\prime},w^{\prime}\in W^{\prime},n^{\prime}\in N^{\prime})$, for an associated evolution function $g^{\prime}=T^{k}$.
With obvious generalizations, we can weaken the RPCT further, by restricting the set of $w^{\prime}\in W^{\prime}$ and / or the set of $\Delta t$ — which corresponds to restrictions on the set of $y$ in Definition 4. Transitivity of simulation (Lemma 4) would still apply for a set of universes related this way. So the simulation graph for such a set of universes would again be a preorder.
@@ 9 Mathematical issues raised by the self-simulation and simulation lemmas
There are also interesting computational complexity issues concerning the three computable functions that define one universe’s simulating another. In particular, there are obvious extensions of the basic framework to concern not perfect simulation, but rather approximate simulation. This then immediately suggests investigating variants of simulation graphs, defined by the approximation complexity of (imperfect) simulation [6]. For example, one might consider simulation graphs where edges from $V$ to $V^{\prime}$ must respect an upper bound on how fast the function ${\mathcal{T}}(\Delta t^{\prime},w^{\prime},n^{\prime})$ would have to grow (as a function of $\Delta t^{\prime}$), in order for the resultant simulation of $V^{\prime}$’s evolution to be at least a factor $\alpha$ within being exactly correct. As another example, one could consider such approximation complexity in the computation of the functions ${\mathcal{T}},{\mathcal{W}},{\mathcal{N}}$ themselves, rather than in the resultant behavior of $\mathcal{T}$}.
Similarly, we can consider the average-case complexity of all the issues arising in the simulation framework. In this kind of approach we would again consider a variant of simulation graphs, this time defined by requiring all edges from one universe (labeled as $V$) to a potentially different universe (labeled as $V^{\prime}$) in the simulation graph must obey associated bounds. In this case though, those bounds would concern the average-case behavior of ${\mathcal{T}}(\Delta t^{\prime},w^{\prime},n^{\prime})$ as a function of $\Delta t^{\prime}$, or of the resources needed to compute the functions ${\mathcal{T}},{\mathcal{W}},{\mathcal{N}}$. We might also want to follow Levin in how precisely to define such “average-case complexity”.
@@ 10 Discussion
Finally, it is worth re-emphasizing that while couched in the language of the simulation hypothesis, the analysis in this paper is far more general. If we assume the Church-Turing thesis, the responses of physical reality during any observation or experiment we might conduct can be modeled as the results of a computation on a TM. The Church-Turing thesis also means that the behavior of we the observers / experimenters can be modeled that way. One might also assume that we ourselves can run a Turing complete computer. This paper is simply an investigation of the formal consequences if we adopt one and / or the other of these two assumptions.
@@ Acknowledgments:
APPENDICES
Appendix A Turing machines
@@
There are many different definitions of Turing machines that are computationally equivalent to one another, in that any computation that can be done with one type of Turing machine can be done with the other. In fact, the “scaling function” of one type of Turing machine, mapping the size of a computation to the minimal amount of resources needed to perform that computation by that type of Turing machine, is at most a polynomial function of the scaling function of any other type of Turing machine. (See for example the relation between the scaling functions of single-tape and multi-tape Turing machines [6].) The following definition will be useful for our purposes, even though it is more complicated than strictly needed:
@@ A Turing machine (TM) is a 7-tuple $(R,\Lambda,b,v,r^{\varnothing},r^{A},\rho)$ where:
Write the set of non-blank symbols of $\Lambda$ as $\hat{\Lambda}$. Every $\lambda$ on the tape of a TM that it might have during a computation in which it halts is a finite string of elements in $\Lambda$ delimited by an infinite string of blanks. Accordingly, wolog we often refer to the state space of the tape as $\Lambda^{*}$, with the trailing infinite string of blanks implicit. Note that $\Lambda^{*}$ is countably infinite, in contrast to $\Lambda^{\infty}$.
@@
We say that a total function $f$ from $(\Lambda\setminus\{b\})^{*}$ to itself is recursive, or (total) computable, if there is a TM with input alphabet $\Lambda$ such that for all $x\in(\Lambda\setminus\{b\})^{*}$, the TM computes $f(x)$. If $f$ is instead a partial function, then we say it is partial recursive (partial computable) if there is a TM with input alphabet $\Lambda$ that computes $f(x)$ for all $x$ for which $f(x)$ is defined, and does not halt for any other $x$. (The reader should be warned that in the literature, the term “computable” is sometimes taken to mean partial computable rather than total computable — and in some articles it is sometimes taken to mean either total computable or partial computable, depending on the context.)
@@
As is standard, I fix some (prefix free, computable, invertible) encoding that maps all finite sets of finite bit strings into a single bit string. Overloading notation, I write that encoding as $\langle.\rangle$, $\langle.,.\rangle$…, depending on the number of strings in the set. Moreover, I require the input to a UTM to be encoded as $\langle k,x\rangle$ if it is emulating TM $T^{k}$ running on input string $x$. However, sometimes for clarity of presentation I will leave the angle brackets implicit, and simply write the UTM $U$ operating on input $\langle k,x\rangle$ as $U(k,x)$. In addition, as shorthand, if $x$ is a vector whose components are all bit strings, I will write the encoded version of all of its components as $\langle x\rangle$.
@@
Formally speaking, the set $\mathbb{B}$$^{*}$ is a Cantor set. A convenient probability measure on this Cantor set, sometimes called the fair-coin measure, is defined so that for any binary string $x$ the set of sequences that begin with $\sigma$ has measure $2^{-|\sigma|}$. Loosely speaking, the fair-coin measure of a prefix TM $T$ is the probability distribution over the strings in $T$’s halting set generated by IID “tossing a coin” to generate those strings, in a Bernoulli process, and then normalizing. [^23] Kraft’s inequality guarantees that since the set of strings in the halting set is a prefix-free set, the sum over all its elements of their probabilities cannot exceed $1$, and so it can be normalized. However, in general that normalization constant is uncomputable, as discussed below. Also, in many contexts we can actually assign arbitrary non-zero probabilities to the strings outside the halting set, as long as the overall distribution is still normalizable. See [39]. So any string $\sigma$ in the halting set has probability $2^{-|\sigma|}/\Omega$ under the fair-coin prior, where $\Omega$ is the normalization constant for the TM in question.
The fair-coin prior provides a simple Bayesian interpretation of Kolmogorov complexity: Under that prior, the Kolmogorov complexity of any string $\sigma$ for any prefix TM $T$ is just (the log of) the maximum over all input strings $\sigma^{\prime}$ of the joint probability of $\sigma^{\prime}$ and $\sigma$. (Strictly speaking, this result is only true up to an additive constant, given by the log of the normalization constant of the fair-coin prior for $T$.)
@@
I end this appendix with some terminological comments and definitions that will be useful in the main text. It is conventional when dealing with Turing machines to implicitly assume some invertible computable map $h(.)$ from $\mathbb{Z}$} into $\Lambda^{*}$. Given such a map $h(.)$, we can exploit it to implicitly assume an additional invertible map taking $\mathbb{Q}$ into $\Lambda^{*}$, e.g., by uniquely expressing any rational number as one product of primes, $a$, divided by a product of different primes, $b$; invertibly mapping those two products of primes into the single integer $2^{a}3^{b}$; and then evaluating $h(2^{a}3^{b})$. Using these definitions, we say that a real number $z$ is computable iff there is a recursive function $f$ mapping rational numbers to rational numbers such that for all rational-valued accuracies $\epsilon>0$, $|f(\epsilon)-z|<\epsilon$. We define computable functions from $\mathbb{Q}\rightarrow{\mathbb{R}}$ similarly.
Appendix B The recursion theorem and Rice’s theorem

### v5

@@ 6.2 Running (self-)simulation using fully homomorphic encryption
Similar consequences arise if we consider self-simulation. We might be simulating ourselves, but have done so with an FHE version of ourselves. Similarl to the possibility discussed above, suppose this is the case, but that we have misplaced the decryption key. In this scenario, we are a simulation that we ourselves are running — but we cannot understand that simulation of ourselves, a simulation which is us.
@@ 7 The simulation graph
Section 6 contains a brief discussion of the philosophical issues that arise if we consider simulation involving more than two universes. There are also interesting mathematical aspects to such a situation. In particular, the graphical structure of universes simulating universes can be quite interesting.
@@ 7.1 The graph of simulations and self-simulations
Indeed, suppose we restrict attention to such a simulation graph universes such that there is a directed edge from any node $V$ in the graph to any other node $V^{\prime}\neq V$. Then the simulation relation cannot be a partially ordered set. Given this, suppose that all of the nodes (universes) have a finite $W$ and an evolution function that obeys the PCT. As pointed out in Section 2.4, the set of such universes is countably infinite, and so we cannot assign a uniform probability distribution to that set. However, since the elements of the set are not partially ordered in the simulation graph, we also cannot assign a Cantor measure over the elements of the set, if we wish to use the simulation relation to fix how to assign such a measure to the nodes in the graph. This establishes the claim made in the introduction, that it is not possible to use a Cantor measure to assign probabilities to a very naturally defined set of universes (at least, it’s not possible if we try to use the simulation relation to fix the measure).
@@
In general, a bare simulation graph could contain computational universes that exist in the same cosmological universe (see Section 2.2), obeying the same laws of physics, the same initial conditions, etc. It might also contain computational universes that exist in different cosmological universes. These two cases can be intermingled as well.
It’s worth briefly commenting on how this second possibility might come about in our particular cosmological universe. One way is if those beings are all running the simulations that are us at the same (cosmological, co-moving) time. In such a situation we would have “split” identities, but at least they would all exist at the same moment in time in our universe. Alternatively though, those beings could be running simulations that are us, but are doing so at different times (i.e., where there is no co-moving frame that contains all of those beings at the moments they are running the simulations). In such a situation, our separate (but identical) selves would all exist at a different moment of time. (Though of course, those versions of us could not have seen that difference in times, at least not yet, since after all, they are identical.)
The simplest way either of these situations could be arranged is if each of the beings running simulations of us exists in a region of the same cosmological universe, where each region is causally disconnected from the regions containing the other beings, and whose backward light cone doesn’t intersect the backward light cones of the other beings simulating us. These restrictions would prevent any complications from the need to have the intersections of the light cones not prevent those beings from all simulating us. (This is true of both the case where all the beings simulating us exist at the same time, and the case where some of them exist at different times.)
In addition to being simulated by multiple other sets of aliens, we could ourselves be running simulations that are us, while perhaps simulating some of those beings who are running simulations of us. In this structure there would be edges that are loops from us into us, and edges from those other universes into us, perhaps together with some edges from us into those other universes.
@@
Suppose we restrict the simulation graph $\Gamma$ to only contain universes that can simulate themselves, Suppose as well that we restrict the graph so that there is at most a single edge from any universe to $V$ to $V^{\prime}$ ( where $V^{\prime}$ may or may not differ from $V$). Then the directed edges in $\Gamma$ form a reflexive, transitive relation, i.e., they form a preorder. In general, the relation provided by the edges in that simulation graph can include both pairs of nodes that are symmetric under the relation and pairs of nodes that are anti-symmetric under the relation. So while it is a preorder, those edges need not provide either an equivalence relation or a partially ordered set.
@@ 7.2 Time-ordered and time-bounded simulation graphs
If we consider a set of computable universes all of which obey the RPCT for $K={\mathbb{N}}$, then the simulation graph is trivial: it is a fully connected graph. However, even if we’re considering a set of computable universes all of which obey RPCT for $K={\mathbb{N}}$, there might still be nontrivial structure in the graph given by placing an edge from $V$ to $V^{\prime}$ only if $V$ simulates $V^{\prime}$ sufficiently quickly. More formally, we can consider the time-bounded simulation graph in which there is an edge from $V$ to $V^{\prime}$ iff $V$ simulates $V^{\prime}$ with a function ${\mathcal{T}}(\Delta t^{\prime},w^{\prime},n^{\prime})$ that obeys some bound in the worst-case over all pairs $(w^{\prime},n^{\prime})$ in how fast it can grow as a function of $\Delta t^{\prime}$. For example, one could consider the variant of a simulation graph given by only placing an edge from $V$ to $V^{\prime}$ if for all pairs $(w^{\prime},n^{\prime})$, ${\mathcal{T}}(\Delta t^{\prime},w^{\prime},n^{\prime})$ grows at most polynomially with $\Delta t^{\prime}$. Other kinds of nontrivial simulation graphs arise if we consider the scaling of the resources (time, memory, etc.) needed to compute the functions ${\mathcal{T}},{\mathcal{W}},{\mathcal{N}}$, or in the case of self-simulation, $\mathcal{S}$}. (See also the discussion in Section 10 of time-minimal simulation functions, their scaling properties, and computational complexity theory.)
@@
Recall as well from Section 4.2 that when the conditions in Lemma 2 hold for a universe $V$. there are an infinite number of initial states of the computer, $n_{0}$, such that that computer simulates the full universe $V$, including itself. This may have some interesting philosophical and mathematical implications. For example, it suggests the possibility for me to run a computer which is simulating my entire universe for a given time into the future $\Delta t$, for my actual environment $w_{0}$, and for a program $n_{0}$ that is computationally equivalent to the actual simulation program I am using — but where that program $n_{0}$ actually differs from the precise program I am using. (So it is not a perfect self-simulation, in that sense.) This suggests replacing any single loop in the simulation graph(i.e., any single edge from a universe-node into itself) with a set of multiple such loops, distinguished by the fact that they use different (but computationally equivalent) programs.
@@ 7.3 Weak RPCT
In general we do not need to assume the full strength of the RPCT to prove a particular instance of either the simulation lemma or the self-simulation lemma. In the case of the former, we just need to assume that the simulating universe can implement the evolution function of the universe being simulated, i.e., can implement the particular TM specified by that evolution function of the universe being simulated. It is not necessary that it can implement the Turing machine of any universe there is. In the case of the latter, we just need to assume that the universe can implement the (Turing machine specifying the) computable function ${\mathcal{S}}(\Delta t)$. We do not need it to be computationally universal.
Accordingly, I say that the weak RPCT (for a set $K\subset{\mathbb{N}}$) holds for universe $V$ if Definition 4 holds for $V$ after one replaces the requirement that the functions $\widehat{{\mathcal{T}}},\widehat{{\mathcal{W}}}$ and $\widehat{{\mathcal{N}}}$ have the RPCT properties for all $k$, instead only requiring they they have the RPCT properties for all $k\in K$ (and evolution function $g$ of $V$).
The associated weak simulation lemma says that $V$ can simulate $V^{\prime}=W^{\prime}\times N^{\prime}$ if $V^{\prime}$ obeys the PCT and $V$ obeys the weak RPCT for a set $K$ that includes three functions $\widehat{{\mathcal{T}}},\widehat{{\mathcal{W}}}$ and $\widehat{{\mathcal{N}}}$ that have the RPCT properties for all $(\Delta t^{\prime},w^{\prime}\in W^{\prime},n^{\prime}\in N^{\prime})$ (for the associated evolution function $g^{\prime}$). With obvious generalizations, we can weaken the RPCT further, by restricting the set of $w^{\prime}\in W^{\prime}$ and / or the set of $\Delta t$ — which corresponds to restrictions on the set of $y$ in Definition 4. Transitivity of simulation (Lemma 4) would still apply for a set of universes related this way. So the simulation graph for such a set of universes would again be a preorder.
@@
7.4 Refinements of simulation graphs
There are many other variants of simulation graphs that might be interesting to explore. As an example, consider some computational universe $V=W\times N$ where $W=\bigtimes_{i\in I}A_{i}$ for some set of spaces $\{A_{i}\}$. Suppose that our computational universe is $V^{\prime}=W^{\prime}\times N^{\prime}$, where $W^{\prime}=\bigtimes{j\in J}A_{j}$ for some subset $J\subset I$. Now suppose that the computer $N$ is running a simulation of us. (For example, there might some species of super-aliens who are a part of the environment $W$ distinct from us, who are running the computer $N$ that way.) In this case there would be two instances of our universe, one given by $V^{\prime}$, and one being simulated in $N$. In this case the simulation graph (as defined above) would only have a single edge, from $V$ to $V^{\prime}$. Yet there would be two instances of us, evolving independently. To complicate the situation further, we might also be running a simulation of ourselves.
In addition, any particular universe might be running more than one simulation at once. Physically, this could occur by having one computer in that universe running multiple simulations simultaneously, just like a laptop runs multiple tasks simultaneously, by “swapping”. It could also occur by having multiple computers in the universe, all running simulations. [^15] In terms of the formalism in this paper, the latter case would mean that “the” computer $N$ of the universe is actually a set of multiple computers running independently, in parallel. The simulation lemma would apply directly, and the self-simulation lemma would also hold, where ${\mathcal{S}}(\Delta t)$ is the initial joint state of all of the computers in the universe.
As an example, there might be a universe $V$ that is running simulations of multiple different universes, $V_{1},V_{2},\ldots$ simultaneously, and as one runs down some of the simulation paths from some subset of those universes, $V_{i_{1}},V_{i_{2}},\ldots$, one eventually converges at us. Similarly, there might be beings in a universe who are running us in their computer who themselves are a simulation in a computer that we are running. In such a case, the aliens would be a simulation running in a computer that they themselves control, just “one step removed”. The same would also be true of us of course. We would be a means for the aliens to “split” their ontological status in two, while also being a means for us to split our ontological status in two.
This suggests an extension of the simulation graph, in which all the edges coming out of each node $V$ are labeled by one or more elements of $\mathbb{N}$}. The interpretation would be that for all $i\in{\mathbb{N}}$, the set of edges out of $V$ labeled $i$ are a maximal set of simulations that $V$ could be doing simultaneously. (“Simultaneous simulation” could be formalized by modifying Definition 1 so that the same initial condition of the simulating computer, $n_{0}\in N$, would result in the computation of the future state of multiple evolution functions.)
@@ 9 Mathematical issues raised by the self-simulation and simulation lemmas
There are also interesting computational complexity issues concerning the three computable functions that define one universe’s simulating another. In particular, there are obvious extensions of the basic framework to concern not perfect simulation, but rather approximate simulation. This then immediately suggests investigating variants of simulation graphs, defined by the the approximation complexity of (imperfect) simulation [5]. For example, one might consider simulation graphs where edges from $V$ to $V^{\prime}$ must respect an upper bound on how fast the function ${\mathcal{T}}(\Delta t^{\prime},w^{\prime},n^{\prime})$ must grow as a function of $\Delta t^{\prime}$ in order for the resultant simulation of $V^{\prime}$’s evolution to be at least a factor $\alpha$ within exact. As another example, one could consider such approximation complexity in the computation of ${\mathcal{T}},{\mathcal{W}},{\mathcal{N}}$ themselves, rather than in the resultant behavior of $\mathcal{T}$}.
Similarly, we can consider the average-case complexity of all the issues arising in the simulation framework. In this kind of approach we would again consider a variant of simulation graphs, this time defined by requiring all edges from one universe (labeled as $V$) to a potentially different universe (labeled as $V^{\prime}$) in the simulation graph must obey associated bounds. In this case though those bounds would concern the average-case behavior of ${\mathcal{T}}(\Delta t^{\prime},w^{\prime},n^{\prime})$ as a function of $\Delta t^{\prime}$, or of the resources needed to compute the functions ${\mathcal{T}},{\mathcal{W}},{\mathcal{N}}$. Why might also want to follow Levin in how precisely to define such “average-case complexity”.
@@ Acknowledgments:
Appendix A Appendix A: Turing machines
@@
There are many different definitions of Turing machines that are computationally equivalent to one another, in that any computation that can be done with one type of Turing machine can be done with the other. It also means that the “scaling function” of one type of Turing machine, mapping the size of a computation to the minimal amount of resources needed to perform that computation by that type of Turing machine, is at most a polynomial function of the scaling function of any other type of Turing machine. (See for example the relation between the scaling functions of single-tape and multi-tape Turing machines [5].) The following definition will be useful for our purposes, even though it is more complicated than strictly needed:
@@ A Turing machine (TM) is a 7-tuple $(R,\Lambda,b,v,r^{\varnothing},r^{A},\rho)$ where:
Write the set of non-blank symbols of $\Lambda$ as $\hat{\Lambda}$. Every $\lambda$ on the tape of a TM that it might have during a computation in which it halts is a finite string of elements in $\Lambda$ delimited by an infinite string of blanks. Accordingly, wolog we often refer to the state space of the tape is $\Lambda^{*}$, with the trailing infinite string of blanks implicit. Note that $\Lambda^{*}$ is countably infinite, in contrast to $\Lambda^{\infty}$.
@@
We say that a total function $f$ from $(\Lambda\setminus\{b\})^{*}$ to itself is recursive, or (total) computable, if there is a TM with input alphabet $\Lambda$ such that for all $x\in(\Lambda\setminus\{b\})^{*}$, the TM computes $f(x)$. If $f$ is instead a partial function, then we say it is partial recursive (partial computable, resp.) if there is a TM with input alphabet $\Lambda$ that computes $f(x)$ for all $x$ for which $f(x)$ is defined, and does not halt for any other $x$. (The reader should be warned that in the literature, the term “computable” is sometimes taken to mean partial computable rather than total computable — and in some articles it is sometimes taken to mean either total computable or partial computable, depending on the context.)
@@
As is standard, I fix some set of (prefix free) encodings of all tuples of finite bit strings, $\langle.\rangle$, $\langle.,.\rangle$… In particular, in general the input to a UTM is encoded as $\langle k,x\rangle$ if it is emulating TM $T^{k}$ running on input string $x$. However, sometimes for clarity of presentation I will leave the angle brackets implicit, and simply write the UTM $U$ operating on input $\langle k,x\rangle$ as $U(k,x)$. In addition, as shorthand, if $x$ is a vector whose components are all bit strings, I will write the encoded version of all of its components as $\langle x\rangle$.
@@
Formally speaking, the set $\mathbb{B}$$^{*}$ is a Cantor set. A convenient probability measure on this Cantor set, sometimes called the fair-coin measure, is defined so that for any binary string $x$ the set of sequences that begin with $\sigma$ has measure $2^{-|\sigma|}$. Loosely speaking, the fair-coin measure of a prefix TM $T$ is the probability distribution over the strings in $T$’s halting set generated by IID “tossing a coin” to generate those strings, in a Bernoulli process, and then normalizing. [^18] Kraft’s inequality guarantees that since the set of strings in the halting set is a prefix-free set, the sum over all its elements of their probabilities cannot exceed $1$, and so it can be normalized. However, in general that normalization constant is uncomputable, as discussed below. Also, in many contexts we can actually assign arbitrary non-zero probabilities to the strings outside the halting set, so long as the overall distribution is still normalizable. See [36]. So any string $\sigma$ in the halting set has probability $2^{-|\sigma|}/\Omega$ under the fair-coin prior, where $\Omega$ is the normalization constant for the TM in question.
The fair-coin prior provides a simple Bayesian interpretation of Kolmogorov complexity: Under that prior, the Kolmogorov complexity of any string $\sigma$ for any prefix TM $T$ is just (the log of) the maximum a posterior (MAP) probability that any string $\sigma^{\prime}$ in the halting set of $T$ was the input to $T$, conditioned on $\sigma$ being the output of that TM. (Strictly speaking, this result is only true up to an additive constant, given by the log of the normalization constant of the fair-coin prior for $T$.)
@@
I end this appendix with some terminological comments and definitions that will be useful in the main text. It is conventional when dealing with Turing machines to implicitly assume some invertible map $h(.)$ from $\mathbb{Z}$} into $\Lambda^{*}$. Given such a map $h(.)$, we can exploit it to implicitly assume an additional invertible map taking $\mathbb{Q}$ into $\Lambda^{*}$, e.g., by uniquely expressing any rational number as one product of primes, $a$, divided by a product of different primes, $b$; invertibly mapping those two products of primes into the single integer $2^{a}3^{b}$; and then evaluating $Rh(2^{a}3^{b})$. Using these definitions, we say that a real number $z$ is computable iff there is a recursive function $f$ mapping rational numbers to rational numbers such that for all rational-valued accuracies $\epsilon>0$, $|f(\epsilon)-z|<\epsilon$. We define computable functions from $\mathbb{Q}\rightarrow{\mathbb{R}}$ similarly.
Appendix B Appendix B: The recursion theorem and Rice’s theorem